OpenStudy (anonymous):

Which of the following is the simplified form ofthe seventh root of x times the seventh root of x times the seventh root of x

OpenStudy (anonymous):

guess it is \[\sqrt[7]{x^3}\]

OpenStudy (anonymous):

simplified? no .. @satellite73

OpenStudy (anonymous):

there is no such mathematical operation as "simplify" it is the lazy math teacher way of saying "give me the answer i want" however, there is nothing you can do to write this in some other form, leave it like that

OpenStudy (anonymous):

i am assuming the initial problem was \[\sqrt[7]{x}\times \sqrt[7]{x}\times \sqrt[7]{x} \] right?

OpenStudy (anonymous):

correct, but the answer you gave me at first was not an option.

OpenStudy (anonymous):

i think the answer is seventh square root of x^3

OpenStudy (anonymous):

oops do they want exponential notation like \[\large x^{\frac{3}{7}}\]?

OpenStudy (anonymous):

not an option either /: these are the options.. \[21\sqrt{x}\]

OpenStudy (anonymous):

or ..

OpenStudy (anonymous):

ok then it is possible that it was not the seventh root, but rather 7 times the square root

OpenStudy (anonymous):

\[X3/21\]

OpenStudy (anonymous):

did it look like this \(\sqrt[7]{x}\) or this \(7\sqrt{x}\)?

OpenStudy (anonymous):

\[7\sqrt{X}\]

OpenStudy (anonymous):

ok that is not the seventh root of \(x\) that is seven times the square root of \(x\)

OpenStudy (anonymous):

so it was \[7\sqrt{x}\times 7\sqrt{x}\times 7\sqrt{x} \]

OpenStudy (anonymous):

so would the answer be .. \[21\sqrt{X} ?\]

OpenStudy (anonymous):

is it this? \[\sqrt[7]{x}^{3}\]

OpenStudy (anonymous):

Yea thats how the problem looks

OpenStudy (anonymous):

and then the answer would be \[343x\sqrt{x}\]

OpenStudy (anonymous):

thats not an option either /:

OpenStudy (anonymous):

can you post a picture?

OpenStudy (anonymous):

maybe it is \[7\sqrt{x}^3\] is that possible?

OpenStudy (anonymous):

\[21\sqrt{x ^{3}}\] ?

OpenStudy (anonymous):

Which of the following is the simplified form of \[7\sqrt{X} TIMES 7\sqrt{X} TIMES 7\sqrt{X}\]

OpenStudy (anonymous):

thats exactly how it looks

OpenStudy (anonymous):

ok well there is no avoiding \(7\times 7\times 7=343\) then

OpenStudy (anonymous):

answer options are ... as following \[21\sqrt{x}\] x 3over7 x 1over7 x 3over21

OpenStudy (anonymous):

ok i get it now

OpenStudy (anonymous):

it is the seventh root of \(x\) it is \[\sqrt[7]{x}\times \sqrt[7]{x}\times \sqrt[7]{x}\] which is in fact equal to \[\sqrt[7]{x^3}\] they want you to write it in exponential notation as \[\large x^{\frac{3}{7}}\]

OpenStudy (anonymous):

the question should read "write the expression using exponential notation"

OpenStudy (anonymous):

thankyouu ! and I agree No idea lets see if i get it right ill let you know i really appreciate your help

OpenStudy (anonymous):

yw is there any way you can post a screen shot of this? is it on line?

OpenStudy (anonymous):

its on my online school its a hw assighment i tried but it doesnt allow me to copy it "/

OpenStudy (anonymous):

ok i was hoping you could just take a screen shot and post that it would be very helpful